OpenStudy (anonymous):

Simplify.. -2 square root p * 2 square root 22

OpenStudy (anonymous):

\[-2\sqrt{p}*2\sqrt{22}=-2*2*\sqrt{p}*22=-4\sqrt{p}*\sqrt{22}=-4\sqrt{22p}\]

OpenStudy (lgbasallote):

\[\LARGE -2 \sqrt p \times 2 \sqrt{22} \implies (-2\times 2) \times (\sqrt{p \times 22})\]

OpenStudy (anonymous):

it should say squaroot of 22 rather than 22